A chair of weight 145 N lies atop a horizontal floor; the floor is not frictionless. You push on the chair with a force of F = 41.0 N directed at an angle of 39.0 ∘ below the horizontal and the chair slides along the floor.

Using Newton's laws, calculate n, the magnitude of the normal force that the floor exerts on the chair.

Express your answer in newtons.

n=. N
Answer #1

Body Diagram Free Flose Fsine

Weight of the chair = W = 145 N

Force with which you push on the chair = F = 41 N

Angle at which the force is applied = \theta = 39o below the horizontal

Normal force on the chair from the floor = N

From the free body diagram in the vertical direction,

N = W + FSin\theta

N = 145 + (41)Sin(39)

N = 184.5 N

Magnitude of the normal force that the floor exerts on the chair = 184.5 N